# Universal bounds on optimization of free energy harvesting

### 20 Dec 2023 v3

@@ 3.2   Deterministic (D) regime
$=-R^{b}_{i^{*}i^{*}}(\ln\alpha-\gamma)\geq K\left(\ln\alpha-\gamma\right)\,,$ (S89)
where we used that $\ln\alpha-\gamma<0$ in the last inequality. For the second term on the right side of (S88),
$\bm{p}^{T}\boldsymbol{\phi}=\sum_{i}p_{i}\phi_{i}=(1-\alpha)\sum_{i}\delta_{ii^{*}}\phi_{i}+\alpha\sum_{i}\pi^{b}_{i}\phi_{i}=(1-\alpha)\mathscr{G}_{\text{D}}+\alpha\dot{\mathcal{G}}({\bm{\pi}^{b}}).$ (S90)
We can plug (3.2) and (S90) into (S88) to give a lower bound on $\mathscr{G}$,
@@
where we used the definition of $\alpha$ (S83) and rearranged. Since $\ln\alpha-\gamma<0$, we can further drop the $\alpha^{2}$ term to give
@@
$\geq\mathscr{G}_{\text{D}}\alpha(\ln\alpha-\gamma-1)+\mathscr{G}_{\text{D}}\,,$ (S91)
@@
$\mathscr{G}_{\text{D}}\alpha\left(\ln\alpha-\gamma-1\right)\leq\mathscr{G}-\mathscr{G}_{\text{D}}\leq\mathscr{G}_{\text{D}}\alpha\gamma\leq\mathscr{G}_{\text{D}}\alpha(-\ln\alpha+\gamma+1)\,,$ (S92)
which can be rearranged to give (S85).
@@
$\left|\frac{\mathscr{G}-\dot{\mathcal{G}}({\bm{\pi}^{b}})}{\mathscr{G}_{\text{D}}-\dot{\mathcal{G}}({\bm{\pi}^{b}})}-1\right|\leq\alpha\left(-\ln\alpha+\gamma+1\right)\,,$ (S93)
@@
3.3   Near-Deterministic (ND) regime
In this regime, we consider a perturbation of the Deterministic regime: we assume that the optimal $\bm{p}^{*}$ is close to delta function distribution centered at $i^{*}={\text{argmax}}_{i}\phi_{i}$. If $p_{i}\approx\delta_{ii^{*}}$,
$\dot{p}_{i}=\sum_{j}R^{b}_{ij}p_{j}\approx R^{b}_{ii^{*}}\,.$ (S94)
This allows us to approximate the entropic term in (S80) as
@@ 3.3   Near-Deterministic (ND) regime
$=R^{b}_{i^{*}i^{*}}\ln\Big{(}1-\sum_{i:i\neq i^{*}}p_{i}\Big{)}+\sum_{i:i\neq i^{*}}R^{b}_{ii^{*}}\ln p_{i}\approx\sum_{i:i\neq i^{*}}\left(-R^{b}_{i^{*}i^{*}}p_{i}+R^{b}_{ii^{*}}\ln p_{i}\right)$ (S95)
where we used $\ln(1-x)\approx-x$ when $x\approx 0$. Plugging into (S80) gives
@@
$=\mathscr{G}_{\text{D}}+\max_{\bm{p}}\sum_{i:i\neq i^{*}}\left[-p_{i}\left(\phi_{i^{*}}-\phi_{i}+\beta^{-1}R^{b}_{i^{*}i^{*}}\right)+\beta^{-1}R^{b}_{ii^{*}}\ln p_{i}\right]=:\mathscr{G}_{\text{ND}}\,,$ (S96)
where we recall that $\mathscr{G}_{\text{D}}=\phi_{i^{*}}$, which does not depend on $\bm{p}$. After maximizing with respect to $\{p_{i}\}_{i:i\neq i^{*}}$ by taking derivatives and setting them to zero, one obtains:
$(\bm{p}^{*}_{\mathrm{ND}})_{i}=\frac{R^{b}_{ii^{*}}}{\beta(\phi_{i^{*}}-\phi_{i})+R^{b}_{i^{*}i^{*}}}\qquad(\text{for}\,i\neq i^{*})\qquad{\text{and}}\qquad\,(\bm{p}^{*}_{\mathrm{ND}})_{i^{*}}=1-\sum_{i:i\neq i^{*}}p_{i}^{*}\,.$ (S97)
Plugging (S97) into (3.3) gives
$\mathscr{G}_{\text{ND}}=\mathscr{G}_{\text{D}}+{\beta}^{-1}\sum_{i:i\neq i^{*}}R^{b}_{ii^{*}}\left[\ln(\bm{p}^{*}_{\mathrm{ND}})_{i}-1\right]\,.$ (S98)
@@
Let us briefly discuss the validity of the expressions above. As stated above, (S97) is valid when $p_{i}\approx\delta_{ii^{*}}$. For convenience, let $i^{-}=\text{arg}\max_{i\neq i^{*}}\phi_{i}$ indicate the second most valuable state of the system after $i^{*}$. We define the harvesting gap as
$\Delta{\phi_{\min}}:=\phi_{i^{*}}-\phi_{i^{-}}=\min_{i:i\neq i^{*}}\phi_{i^{*}}-\phi_{i}\,.$
Now, considering the optimal probability for $i\neq i^{*}$ found in (S97), the ND approximation will be valid as long as
$p^{*}_{i\neq i^{*}}\ll 1\quad\Leftrightarrow\quad\phi_{i^{*}}-\phi_{i}\gg{\beta}^{-1}\left(R^{b}_{ii^{*}}-R^{b}_{i^{*}i^{*}}\right)\,,$
which holds as long as
$\Delta{\phi_{\min}}\gg{\beta}^{-1}\left(R^{b}_{i^{-}i^{*}}-R^{b}_{i^{*}i^{*}}\right)\,.$ (S99)
Expression (S99) implies that ND approximation is valid when the harvesting gap is much larger than the rates at which either $i^{*}$ flows into $i^{-}$ or escapes.
Unicyclic model
@@
4.1   An Algebraic Aperitif: eigendecomposition of unicyclic rate matrix
@@
$R^{b}=\left(\begin{matrix}-2k&k&0&\cdots&k\\ k&-2k&k&\cdots&0\\ 0&k&-2k&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ k&0&0&\cdots&-2k\end{matrix}\right)\,,$ (S105)
which is a symmetric circulant matrix. Due to symmetry, its steady state is uniform: $\pi^{b}_{i}=\frac{1}{n}$, $\forall i$. The eigensystem for $R$ is obtained from the theory of circulant matrices [55], which, for odd $n$, yields:
$\lambda_{a}$ $=-2k\left[1-\Re\left(\omega_{a}\right)\right]=-2k\left[1-\cos\left(\frac{2\pi(a-1)}{n}\right)\right]\,,$ (S106)
@@
$\{m_{(a)}\}=\left\{\frac{1}{\sqrt{n}}\left(1,\omega_{a},\omega_{a}^{2},\ldots,\omega_{a}^{n-1}\right)\right\}.$ (S107)
4.2   Transition harvesting cycle
We now analyze the unicyclic model using techniques developed in Sec. 3. We consider a systems with $n$ state arranged in a ring, where baseline transitions are symmetric with uniform kinetics: $i\stackrel{{\scriptstyle[}}{{k}}]{k}{\leftrightharpoons}i+1\mod n$, $\forall i=1,\ldots,n$ (see Fig. S11 here and Fig. 2(a) in the main text). The baseline dynamics are equivalent to a random walk on a one-dimensional ring, and the baseline rate matrix is given by (S105).
@@ 4.2   Transition harvesting cycle
$\boldsymbol{\phi}^{\text{LR}}=\boldsymbol{\phi}=(k\Theta,-k\Theta,0,\ldots,0)\,.$ (S108)
@@
4.2.1   LR regime
We first consider the LR solution (S77). Combining (S108) with (S72), $\pi^{b}_{i}=1/n$, and $D=\frac{1}{\sqrt{n}}I_{n}$ gives $\boldsymbol{v}=\frac{k\beta\Theta}{2\sqrt{n}}(1,-1,0,\ldots,0)$. In addition, since in this particular example $R$ obeys detailed balance, $R^{b}=A$. Finally, it is easy to verify that $M=D^{-1}AD=A=R^{b}$. Thus, $M$ has the same eigendecomposition as $R$. The normalized eigenvectors of $M$ are given by (S107).
We can compute $\Omega_{a}={\beta}^{-1}\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}$ for $a=2,\ldots,n$, using the eigenvector set given in (S107),which simply yields
$\Omega_{a}=\frac{k\Theta}{2n}\left(1-\omega_{a}\right)\,.$ (S109)
Next, we compute the maximum harvesting rate attainable using (S77),
@@ 4.2.1   LR regime
$=\beta k\left(\frac{\Theta}{2n}\right)^{2}\sum_{a=2}^{n}1=\beta k\left(\frac{\Theta}{2n}\right)^{2}(n-1)\,.$ (S110)
On the other hand, continuing with odd $n$, it is also possible to compute the deviation of the optimal distribution from stationary distribution using (S77),
$\Delta\bm{p}^{*}_{\mathrm{LR}}:=\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}=\beta\sum_{a=2}^{n}\frac{\Omega_{a}^{\dagger}D\boldsymbol{m}^{\alpha}}{-\lambda_{a}}\,,$ (S111)
which is depicted in Fig. S12(a). (Note that, due to our choice of a complex-valued eigenbasis, we must be careful in adding the complex conjugate on the $\Omega_{a}$, following the ordering of the operator $M^{+}$ in (S77).) Now, component by component, we can rewrite (S111) as
@@
$=\frac{\beta\Theta}{4n^{2}}\left[-\sum_{a=2}^{n}\cos\left(\frac{2\pi(a-1)(i-2)}{n}\right)-\sum_{a=2}^{n}\frac{\sin\left(\frac{2\pi(a-1)(i-2)}{n}\right)\sin\left(\frac{2\pi(a-1)}{n}\right)}{1-\cos\left(\frac{2\pi(a-1)}{n}\right)}\right].$ (S112)
In the first line, we used the expressions for $\Omega_{a}$ from (S109), $\lambda_{a}$ from (S106), $\boldsymbol{m}^{\alpha}$ from (S107), and $D=I/\sqrt{n}$ and then simplified. In the second line, we expanded $(1-\omega_{a})^{\dagger}\omega_{a}^{i-1}=\omega_{a}^{i-1}-\omega_{a}^{i-2}$ into real and imaginary components and then used that the imaginary components cancel over the sum. In the last line, we used the identity $\cos(x+y)=\cos(x)\cos(y)-\sin(x)\sin(y)$ for $x=\frac{2\pi(a-1)(i-2)}{n}$ and $y=\frac{2\pi(a-1)}{n}$ and then simplified. It can be verified that the first series in (S112) sums to $-1$. The second series is trickier but can be simplified using trigonometric methods as shown in Ref. [56]. Plugging in that solution and simplifying gives the very simple expression:
$\left(\Delta\bm{p}^{*}_{\mathrm{LR}}\right)_{i}=\frac{\beta\Theta}{4n^{2}}\begin{cases}2(i-1)-(n+1)&{\text{\qquad for }}i=2,\ldots,n\\ (n-1)&{\text{\qquad for }}i=1\end{cases}\,.$ (S113)
@@
Figure S12: The optimal $\Delta\bm{p}^{*}_{\mathrm{LR}}$ (deviation of optimal distribution from the uniform baseline steady state) given (S113). The optimal distribution exhibits a clockwise cyclic current given by (S114), The magnitude of $\Delta\bm{p}^{*}_{\mathrm{LR}}$ decays as $\sim n^{-1}$, so larger rings will result in optimal distributions that are closer to the baseline steady state.
@@
$k{\left[(\bm{p}^{*}_{\mathrm{LR}})_{1}-(\bm{p}^{*}_{\mathrm{LR}})_{2}\right]}=\frac{k\beta\Theta}{2n}\left(\frac{n-1}{n}\right)\,.$ (S114)
The expressions (S78) and (S79) suggest for which values of $\Theta$ the optimal solution will be in the LR regime. In this case, however, it is possible to work out the norm $\|\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}\|$ exactly:
@@
$\|\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}\|\ll 1\Leftrightarrow\Theta\ll\frac{4\sqrt{3}}{\beta}\sqrt{\frac{n^{3}}{n^{2}-1}}$ (S115)
@@
4.2.2   D and ND Regimes
Note that the optimal mesostate is $i^{*}=1$. Hence $\mathscr{G}_{\text{D}}=\phi_{1}=k\Theta$ and $(\bm{p}^{*}_{\mathrm{D}})_{i}=\delta_{i1}$. Next, consider the ND expressions (S97) with (S108). Denote $\Delta\phi_{2}=\phi_{i^{*}}-\phi_{2}=2k\Theta$ and $\Delta\phi_{n}=\phi_{i^{*}}-\phi_{n}=k\Theta$; and observe that $R^{b}_{21}=R^{b}_{n1}=k$, $R^{b}_{11}=-2k$ and $R^{b}_{j1}=0$ for all $j\neq 2,n$. Hence, for $i\neq i^{*}$, we obtain
$(\bm{p}^{*}_{\mathrm{ND}})_{i}&=\frac{R^{b}_{i1}}{\beta\Delta\phi_{i}+R^{b}_{11}}=\begin{cases}\frac{1}{2(\beta\Theta-1)}&\qquad i=2\\ \frac{1}{\beta\Theta-2}&\qquad i=n\\ 0&\qquad{\text{otherwise}}\end{cases}\\ \left(\bm{p}^{*}_{\mathrm{ND}}\right)_{i^{*}=1}&=1-\sum_{i:i\neq i^{*}}(\bm{p}^{*}_{\mathrm{ND}})_{i}=1-\frac{3\beta\Theta-4}{2(\beta\Theta-1)(\beta\Theta-2)}\,.\end{aligned}$ (S116)
The free energy harvesting rate attainable in this regime is given by
$\mathscr{G}_{\text{ND}}$ $=\phi_{1}+{\beta^{-1}}\sum_{i:i\neq i^{*}}R^{b}_{i1}\left[\ln\left(\bm{p}^{*}_{\mathrm{ND}}\right)_{i}-1\right]$ (S117)
$=k(\Theta-2/\beta)-k\beta^{-1}\ln\left[2(\beta\Theta-1)(\beta\Theta-2)\right].$ (S118)
4.3   State harvesting cycle
In this section, we consider an alternative unicyclic model (see Fig. S13), which is different from the one studied in the main text and in Sec. 4.
In this model, the baseline dynamics correspond once again to the unicycle with symmetric rates, as in Sec. 4.2. However, harvesting is not coupled to transitions, but rather to internal fluxes within a single mesostate. Without loss of generality, we choose this mesostate to be $i=1$. In other words, we let $\dot{g}^{b}_{1}=\theta>0$ and $\dot{g}^{b}_{1<i\leq n}=0$. There is now no preferential direction for the probability current. The free energy harvested per unit time when the system is in mesostate $1$ is given by the parameter $\theta$, which carries the same units as $\mathscr{G}$. Just as before, the baseline steady state is uniform, i.e. $\pi^{b}_{i}=1/n$. Hence, the harvesting vector is
$\boldsymbol{\phi}^{\text{LR}}=\boldsymbol{\phi}=(\theta,0,\ldots,0)\,.$ (S119)
This section provides a complementary illustration of how our theoretical framework unfolds in a different scenario. Note that this model is not discussed in the main text, and only appears here in the SM. Below we consider both the LR, D and ND regimes.

### 12 Mar 2024 v4

@@ II.4 Proton-pumping as control
$=-R^{b}_{i^{*}i^{*}}(\ln\alpha-\gamma)\geq K\left(\ln\alpha-\gamma\right)\,,$ (S3.88)
where we used that $\ln\alpha-\gamma<0$ in the last inequality. For the second term on the right side of (S3.87),
$\boldsymbol{\phi}^{T}\bm{p}=\sum_{i}p_{i}\phi_{i}=(1-\alpha)\sum_{i}\delta_{ii^{*}}\phi_{i}+\alpha\sum_{i}\pi^{b}_{i}\phi_{i}=(1-\alpha)\mathscr{G}_{\text{D}}+\alpha\dot{\mathcal{G}}({\bm{\pi}^{b}}).$ (S3.89)
We can plug (III.2) and (S3.89) into (S3.87) to give a lower bound on $\mathscr{G}$,
@@
where we used the definition of $\alpha$ (S3.82) and rearranged. Since $\ln\alpha-\gamma<0$, we can further drop the $\alpha^{2}$ term to give
@@
$\geq\mathscr{G}_{\text{D}}\alpha(\ln\alpha-\gamma-1)+\mathscr{G}_{\text{D}}\,,$ (S3.90)
@@
$\mathscr{G}_{\text{D}}\alpha\left(\ln\alpha-\gamma-1\right)\leq\mathscr{G}-\mathscr{G}_{\text{D}}\leq\mathscr{G}_{\text{D}}\alpha\gamma\leq\mathscr{G}_{\text{D}}\alpha(-\ln\alpha+\gamma+1)\,,$ (S3.91)
which can be rearranged to give (S3.84).
@@
$\left|\frac{\mathscr{G}-\dot{\mathcal{G}}({\bm{\pi}^{b}})}{\mathscr{G}_{\text{D}}-\dot{\mathcal{G}}({\bm{\pi}^{b}})}-1\right|\leq\alpha\left(-\ln\alpha+\gamma+1\right)\,,$ (S3.92)
@@
III.3 Near-Deterministic (ND) regime, Eq. (17)
In this regime, we consider a perturbation of the Deterministic regime: we assume that the optimal $\bm{p}^{*}$ is close to a delta-function distribution centered at $i^{*}={\text{argmax}}_{i}\phi_{i}$. If $p_{i}\approx\delta_{ii^{*}}$,
$\dot{p}_{i}=\sum_{j}R^{b}_{ij}p_{j}\approx R^{b}_{ii^{*}}\,.$ (S3.93)
This allows us to approximate the entropic term in (S3.79) as
@@
$=R^{b}_{i^{*}i^{*}}\ln\Big{(}1-\sum_{i:i\neq i^{*}}p_{i}\Big{)}+\sum_{i:i\neq i^{*}}R^{b}_{ii^{*}}\ln p_{i}$
$\approx\sum_{i:i\neq i^{*}}\left(-R^{b}_{i^{*}i^{*}}p_{i}+R^{b}_{ii^{*}}\ln p_{i}\right)$ (S3.94)
where we used $\ln(1-x)\approx-x$ when $x\approx 0$. Plugging into (S3.79) gives
@@
$=\mathscr{G}_{\text{D}}+\max_{\bm{p}}\sum_{i:i\neq i^{*}}\left[-p_{i}\left(\phi_{i^{*}}-\phi_{i}+\beta^{-1}R^{b}_{i^{*}i^{*}}\right)+\beta^{-1}R^{b}_{ii^{*}}\ln p_{i}\right]=:\mathscr{G}_{\text{ND}}\,,$ (S3.95)
where we recall that $\mathscr{G}_{\text{D}}=\phi_{i^{*}}$, which does not depend on $\bm{p}$. After maximizing with respect to $\{p_{i}\}_{i:i\neq i^{*}}$ by taking derivatives and setting them to zero, one obtains
$(\bm{p}^{*}_{\mathrm{ND}})_{i}=\frac{R^{b}_{ii^{*}}}{\beta(\phi_{i^{*}}-\phi_{i})+R^{b}_{i^{*}i^{*}}}\qquad(\text{for}\,i\neq i^{*})\qquad{\text{and}}\qquad\,(\bm{p}^{*}_{\mathrm{ND}})_{i^{*}}=1-\sum_{i:i\neq i^{*}}(\bm{p}^{*}_{\mathrm{ND}})_{i}\,.$ (S3.96)
Plugging (S3.96) into (III.3) gives
$\mathscr{G}_{\text{ND}}=\mathscr{G}_{\text{D}}+{\beta}^{-1}\sum_{i:i\neq i^{*}}R^{b}_{ii^{*}}\left[\ln(\bm{p}^{*}_{\mathrm{ND}})_{i}-1\right]\,.$ (S3.97)
@@
The ND approximation applies when $(\bm{p}^{*}_{\mathrm{ND}})_{i^{*}}\approx 1$, or equivalently when
$1\gg\sum_{i:i\neq i^{*}}(\bm{p}^{*}_{\mathrm{ND}})_{i}=\sum_{i:i\neq i^{*}}\frac{R^{b}_{ii^{*}}}{\beta(\phi_{i^{*}}-\phi_{i})+R^{b}_{i^{*}i^{*}}}\,.$
For convenience, we define the harvesting gap as the difference in the pay-off between the optimal state $i^{*}$ and the second optimal state,
$\Delta{\phi_{\min}}:=\phi_{i^{*}}-\max_{i:i\neq i^{*}}\phi_{i}\,.$
By combining and rearranging the above, we can simplify the validity condition for the ND approximation as
$\beta\Delta{\phi_{\min}}\gg-2R^{b}_{i^{*}i^{*}},$ (S3.98)
Note that this condition also guarantees that $(\bm{p}^{*}_{\mathrm{ND}})_{i}\geq 0$ for $i\neq i^{*}$. Expression (S3.98) implies that ND approximation is valid when the harvesting gap is much larger than the rates of escape from the optimal state $i^{*}$.
IV Unicyclic model
@@
IV.1 An Algebraic Aperitif: eigendecomposition of unicyclic rate matrix
@@
$R^{b}=\left(\begin{matrix}-2k&k&0&\cdots&k\\ k&-2k&k&\cdots&0\\ 0&k&-2k&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ k&0&0&\cdots&-2k\end{matrix}\right)\,,$ (S4.104)
which is a symmetric circulant matrix. Due to symmetry, its steady state is uniform: $\pi^{b}_{i}=1/n$ for all $i$. The eigensystem for $R$ is obtained from the theory of circulant matrices [45], which, for odd $n$, yields:
$\lambda_{a}$ $=-2k\left[1-\Re\left(\omega_{a}\right)\right]=-2k\left[1-\cos\left(\frac{2\pi(a-1)}{n}\right)\right]\,,$ (S4.105)
@@
$\{m_{(a)}\}=\left\{\frac{1}{\sqrt{n}}\left(1,\omega_{a},\omega_{a}^{2},\ldots,\omega_{a}^{n-1}\right)\right\}.$ (S4.106)
IV.2 Transition harvesting cycle
[…]
@@ IV.2 Transition harvesting cycle
$\boldsymbol{\phi}^{\text{LR}}=\boldsymbol{\phi}=(k\Theta,-k\Theta,0,\ldots,0)\,.$ (S4.107)
@@
IV.2.1 LR regime
We first consider the LR solution (S3.76). Combining (S4.107) with (S3.71), $\pi^{b}_{i}=1/n$, and $D=\frac{1}{\sqrt{n}}I_{n}$ gives $\boldsymbol{v}=\frac{k\beta\Theta}{2\sqrt{n}}(1,-1,0,\ldots,0)$. In addition, since in this particular example $R$ obeys detailed balance, $R^{b}=A$. Finally, it is easy to verify that $M=D^{-1}AD=A=R^{b}$. Thus, $M$ has the same eigendecomposition as $R$. The normalized eigenvectors of $M$ are given by (S4.106).
We can compute $\Omega_{a}={\beta}^{-1}\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}$ for $a=2,\ldots,n$, using the eigenvector set given in (S4.106),which simply yields
$\Omega_{a}=\frac{k\Theta}{2n}\left(1-\omega_{a}\right)\,.$ (S4.108)
Next, we compute the maximum harvesting rate attainable using (S3.76),
@@ IV.2.1 LR regime
$=\beta k\left(\frac{\Theta}{2n}\right)^{2}\sum_{a=2}^{n}1$
$=\beta k\left(\frac{\Theta}{2n}\right)^{2}(n-1)\,.$ (S4.109)
On the other hand, continuing with odd $n$, it is also possible to compute the deviation of the optimal distribution from stationary distribution using (S3.76),
$\Delta\bm{p}^{*}_{\mathrm{LR}}:=\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}=\beta\sum_{a=2}^{n}\frac{\Omega_{a}^{\dagger}D\boldsymbol{m}^{\alpha}}{-\lambda_{a}}\,,$ (S4.110)
which is depicted in Fig. S14(a). (Note that, due to our choice of a complex-valued eigenbasis, we must be careful in adding the complex conjugate on the $\Omega_{a}$, following the ordering of the operator $M^{+}$ in (S3.76).) Now, component by component, we can rewrite (S4.110) as
@@
$=\frac{\beta\Theta}{4n^{2}}\left[-\sum_{a=2}^{n}\cos\left(\frac{2\pi(a-1)(i-2)}{n}\right)-\sum_{a=2}^{n}\frac{\sin\left(\frac{2\pi(a-1)(i-2)}{n}\right)\sin\left(\frac{2\pi(a-1)}{n}\right)}{1-\cos\left(\frac{2\pi(a-1)}{n}\right)}\right].$ (S4.111)
In the first line, we used the expressions for $\Omega_{a}$ from (S4.108), $\lambda_{a}$ from (S4.105), $\boldsymbol{m}^{\alpha}$ from (S4.106), and $D=I/\sqrt{n}$ and then simplified. In the second line, we expanded $(1-\omega_{a})^{\dagger}\omega_{a}^{i-1}=\omega_{a}^{i-1}-\omega_{a}^{i-2}$ into real and imaginary components and then used that the imaginary components cancel over the sum. In the last line, we used the identity $\cos(x+y)=\cos(x)\cos(y)-\sin(x)\sin(y)$ for $x=\frac{2\pi(a-1)(i-2)}{n}$ and $y=\frac{2\pi(a-1)}{n}$ and then simplified. It can be verified that the first series in (S4.111) sums to $-1$. The second series is trickier but can be simplified using trigonometric methods as shown in Ref. [79]. Plugging in that solution and simplifying gives the very simple expression:
$\left(\Delta\bm{p}^{*}_{\mathrm{LR}}\right)_{i}=\frac{\beta\Theta}{4n^{2}}\begin{cases}2(i-1)-(n+1)&{\text{\qquad for }}i=2,\ldots,n\\ (n-1)&{\text{\qquad for }}i=1\end{cases}\,.$ (S4.112)
@@
Figure S14: The optimal $\Delta\bm{p}^{*}_{\mathrm{LR}}$ (deviation of optimal distribution from the uniform baseline steady state) given (S4.112). The optimal distribution exhibits a clockwise cyclic current given by (S4.113), The magnitude of $\Delta\bm{p}^{*}_{\mathrm{LR}}$ decays as $\sim n^{-1}$, so larger rings will result in optimal distributions that are closer to the baseline steady state.
@@
$k{\left[(\bm{p}^{*}_{\mathrm{LR}})_{1}-(\bm{p}^{*}_{\mathrm{LR}})_{2}\right]}=\frac{k\beta\Theta}{2n}\left(\frac{n-1}{n}\right)\,.$ (S4.113)
The expressions (S3.77) and (S3.78) suggest for which values of $\Theta$ the optimal solution will be in the LR regime. In this case, however, it is possible to work out the norm $\|\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}\|$ exactly:
@@
$\|\bm{p}^{*}_{\mathrm{LR}}-{\bm{\pi}^{b}}\|\ll 1\Leftrightarrow\Theta\ll\frac{4\sqrt{3}}{\beta}\sqrt{\frac{n^{3}}{n^{2}-1}}$ (S4.114)
@@
IV.2.2 D and ND Regimes
Note that the optimal mesostate is $i^{*}=1$. Hence $\mathscr{G}_{\text{D}}=\phi_{1}=k\Theta$ and $(\bm{p}^{*}_{\mathrm{D}})_{i}=\delta_{i1}$. Next, consider the ND expressions (S3.96) with the payoff vector (S4.107). Denote $\Delta\phi_{2}=\phi_{i^{*}}-\phi_{2}=2k\Theta$ and $\Delta\phi_{n}=\phi_{i^{*}}-\phi_{n}=k\Theta$; and observe that $R^{b}_{21}=R^{b}_{n1}=k$, $R^{b}_{11}=-2k$ and $R^{b}_{j1}=0$ for all $j\neq 2,n$. Hence, for $i\neq i^{*}$, we obtain
$(\bm{p}^{*}_{\mathrm{ND}})_{i}&=\frac{R^{b}_{i1}}{\beta\Delta\phi_{i}+R^{b}_{11}}=\begin{cases}\frac{1}{2(\beta\Theta-1)}&\qquad i=2\\ \frac{1}{\beta\Theta-2}&\qquad i=n\\ 0&\qquad{\text{otherwise}}\end{cases}\\ \left(\bm{p}^{*}_{\mathrm{ND}}\right)_{i^{*}=1}&=1-\sum_{i:i\neq i^{*}}(\bm{p}^{*}_{\mathrm{ND}})_{i}=1-\frac{3\beta\Theta-4}{2(\beta\Theta-1)(\beta\Theta-2)}\,.\end{aligned}$ (S4.115)
The harvesting rate attainable in this regime can be calculated using Eq. (S3.97),
$\mathscr{G}_{\text{ND}}$ $=\phi_{1}+{\beta^{-1}}\sum_{i:i\neq i^{*}}R^{b}_{i1}\left[\ln\left(\bm{p}^{*}_{\mathrm{ND}}\right)_{i}-1\right]$ (S4.116)
$=k(\Theta-2/\beta)-k\beta^{-1}\ln\left[2(\beta\Theta-1)(\beta\Theta-2)\right].$ (S4.117)
IV.3 State harvesting cycle
In this section, we consider the second unicyclic model, see Fig. S15 and Fig. 4 (b) in the main text. In this model, the baseline dynamics correspond once again to the unicycle with symmetric rates, as in Sec. IV.2. However, harvesting is not coupled to transitions, but rather to internal fluxes within a single mesostate. Without loss of generality, we choose this mesostate to be $i=1$. In other words, we let $\dot{g}^{b}_{1}=\theta>0$ and $\dot{g}^{b}_{1<i\leq n}=0$. There is now no preferential direction for the probability current. The free energy harvested per unit time when the system is in mesostate $1$ is given by the parameter $\theta$, which carries the same units as $\mathscr{G}$. Just as before, the baseline steady state is uniform, i.e. $\pi^{b}_{i}=1/n$. Hence, the harvesting vector is
$\boldsymbol{\phi}^{\text{LR}}=\boldsymbol{\phi}=(\theta,0,\ldots,0)\,.$ (S4.118)
Below we consider the LR, D and ND regimes.